(Energy problem from last week). During a dunk, a basketball player will elevate their center of mass 70 cm (or more) while jumping off of one leg. What is the strain energy contribution to the energy required to elevate the center of mass given the following parameters? How much energy has to be contributed by active muscle contraction? mass = 82 kg, Kleg = 35000N/m, Δx = 0.11 m.
Answer #1

As footballer elevates his Center of mass by 0.7m and his mass is 82kg then energy required for that action is given by

mgh = 82 × 0.7 × 9.81 = 563.09 j ans

And as the active muscle contracted by 0.11m and Kleg=35000N/m so energy provided by active muscle is given by 1/2 × Kleg × x²

= 1/2 × 35000 × 0.11² = 211.75 j ans